# Steering LLMs for Formal Theorem Proving

### v4

@@ Abstract
Recent progress in automated theorem proving leverages Large Language Models (LLMs) for their capacity to comprehend informal mathematical statements and generate corresponding formal proofs. Even though these techniques perform well, very little exploration has been done to understand how language models interpret and utilize these informal mathematical cues to generate formal proofs more effectively. To address this, we explore activation steering, a lightweight, inference-time mechanism that identifies linear directions in a model’s residual activations corresponding to informal “thought” traces, and nudges those activations to improve proof construction entirely without finetuning. Unlike previous approaches, activation engineering offers valuable insights into language models’ internal reasoning dynamics encoded in their activation space. We evaluated these activation vectors on two distinct tasks: formal proof generation from formal theorems and formal proof generation from informal problem descriptions. Our contributions are twofold: (1) we purpose an activation-based intervention technique to guide proof synthesis in LLMs; and (2) improve performance across two different decoding strategies without additional training.
Machine Learning, ICML
@@ 1 Introduction
Interactive proof assistants such as Lean (de Moura et al.,, 2015), Isabelle (Wenzel et al.,, 2008), and Coq (Barras et al.,, 1999) enable formal verification of mathematical proofs and software by leveraging specialized programming languages (Avigad,, 2023; Ringer et al.,, 2019). Neural theorem proving, which integrates neural models like Large Language Models (LLMs) with these assistants, has emerged as a promising approach to automating formal reasoning and theorem proving (First et al.,, 2023; Polu and Sutskever,, 2020; Polu et al.,, 2022; Yang et al.,, 2023; Welleck,, 2023). This integration is intuitive and powerful: proof assistants enforce formal correctness, while neural models predict and suggest logical steps for proof construction. A key challenge, however, is generalizing proof writing to scenarios unseen during training. Multiple prior studies have demonstrated that informal (natural language, NL) information can enhance theorem proving in language models (Lin et al.,, 2024; Welleck et al.,, 2021, 2022).
Despite these advances, it remains unclear how natural language information influences a model’s reasoning process and whether it can be systematically leveraged to improve proof-writing capabilities. Our work addresses this gap by investigating the mechanistic role of natural language information in LLM reasoning and developing a method to harness it effectively.
We find that the task of generating natural language mathematical context can be encoded as activation vectors in the model’s activation space. Using activation steering (Panickssery et al., 2024b, ; Turner et al.,, 2024; Lucchetti and Guha,, 2024), we capture differences in model behavior driven by natural language inputs, enabling targeted interventions that enhance proof-writing without affecting unrelated aspects of performance. In figure 1 we demonstrate how structured natural language interventions improve the quality and the structure of proof. In comparison to natural sampling leads to counterfeit and incorrect proofs. We evaluate this approach on three open-source LLMs like Lemma (Azerbayev et al.,, 2024), InternLM-2 (Ying et al.,, 2024), and InternLM2.5-StepProver (Wu et al.,, 2024) that are designed for theorem proving through training on mathematics data. Our experiments show that steering vectors with targeted NL information consistently improve proof quality across the MiniF2F and PutnamBench benchmarks.
Our key contributions are:
@@
    Mechanistic insights into how informal mathematical context and natural language information affects LLM reasoning about formal theorems.
@@
    Activation vector identification for encoding natural language information to guide formal proof construction in a structured and grounded manner.

### v5

# Steering Language Models for Theorem Proving

@@ Abstract
Recent advances in automated theorem proving use Large Language Models (LLMs) to translate informal mathematical statements into formal proofs. However, informal cues are often ambiguous or lack strict logical structure, making it hard for models to interpret them precisely. While existing methods achieve strong performance, little is known about how LLMs internally represent informal cues, or how these influence proof generation. To address this, we explore activation steering, an inference-time intervention that identifies linear directions in residual activations associated with informal reasoning traces and adjusts them to improve proof construction without fine-tuning. This mechanism also yields interpretable information about how reasoning is internally encoded in the activation space of LLMs. We test our method for generating formal proofs from already-formalized theorems. Our contributions are twofold: (1) a novel activation-based intervention for guiding proof synthesis in LLMs; and (2) demonstration that this intervention improves performance under two decoding strategies (sampling and best-first search) without any further training.
@@ 1 Introduction
Interactive proof assistants such as Lean (de Moura et al., 2015), Isabelle (Wenzel et al., 2008), and Rocq (Barras et al., 1999) provide the infrastructure for formal verification of mathematical proofs and software. They require proofs to be expressed in precise formal languages (Avigad, 2023; Ringer et al., 2019). Neural theorem proving, combining LLMs with proof assistants, has shown promise in automating reasoning (First et al., 2023; Polu & Sutskever, 2020; Polu et al., 2022; Yang et al., 2023; Welleck, 2023). Prior work (Lin et al., 2024; Welleck et al., 2021; 2022) suggests that training model on natural reasoning for proof steps can improve performance, but it remains unclear how such informal language is internally represented or whether it can be reliably leveraged to improve proof generation.
Building on the observation that proofs often interleave natural-language reasoning with formal steps, we show that informal natural language (NL) context induces distinct activation patterns in LLMs. We extract these patterns as steering vectors as in (Panickssery et al., 2024b; Turner et al., 2024; Lucchetti & Guha, 2024), and use them to intervene at inference time. These interventions improve proof generation quality, while preserving model behavior in unrelated aspects. We test our approach across three theorem-proving oriented LLMs - Lemma (Azerbayev et al., 2024), InternLM-2 (Ying et al., 2024), and InternLM2.5-StepProver (Wu et al., 2024). On established benchmarks MiniF2F and PutnamBench, we find that steering via informal language context consistently improves proof quality under multiple decoding strategies.
Our key contributions are threefold: (1) We provide mechanistic insights into how informal mathematical or natural language context impacts internal reasoning in LLMs for theorem proving. (2) We propose a method to extract activation vectors that encode such informal context, and use these vectors to guide proof construction in a structured, grounded way. (3) We demonstrate that activation steering yields consistent improvements in proof success rates on MiniF2F and PutnamBench benchmarks under both search- and sampling-based decoding.
By focusing on steering in activation space, our approach sheds light on the link between informal mathematical language and formal reasoning steps; without needing to fine-tune model weights. We base this on the hypothesis that many reasoning features are encoded as (approximately) linear directions in activation space (Mikolov et al., 2013; Elhage et al., 2021; Nanda et al., 2023; Park et al., 2024). Although not all features follow perfect linearity (Engels et al., 2025), this assumption has previously enabled methods like concept erasure and steering to work well (Beaglehole et al., 2025; Zhao et al., 2025a; Shah et al., 2025).
The rest of this paper is organized as follows. In Section 2, we review related work on neural theorem proving, activation steering, and the representation of informal mathematical reasoning in LLMs. Section 3 presents our activation steering method: how steering vectors are constructed from informal mathematical contexts, how they are applied at inference time, and how we select layers and intervention strengths. Section 5 describes our experimental setup, including models, benchmarks (MiniF2F, PutnamBench), and decoding strategies (sampling vs best-first search). Section 6 reports results: we analyze performance gains, the effect of steering vectors on proof success rates, and ablations experiments. Finally, in Section 7 we discuss insights into the internal reasoning of the model, limitations of our approach, and potential directions for future work.
2 Related Work
This section situates our work in three intersecting strands: automatic theorem proving and auto‐formalization, representation learning in large language models (LLMs), and mechanistic interpretability via activation interventions.
Automatic Theorem Proving and Auto-formalization Recent progress in automatic theorem proving frames proof search as sequence generation. The GPT-f framework (Polu & Sutskever, 2020) trains a language model to map proof states to tactics, combining it with best‐first search to assemble proofs. Extensions explore data augmentation via proof transformations or synthesis (Han et al., 2022; Rotella et al., 2025; Wang & Deng, 2020), improved search strategies (Wang et al., 2023), curriculum training (Polu et al., 2022), and retrieval from proof libraries (Yang et al., 2023). Systems like LLMStep unify these ideas into usable frameworks (Welleck & Saha, 2023).
Auto‐formalization complements this by translating informal mathematics (e.g. text, sketches) into formal proofs. Surveys highlight translation techniques (Wu et al., 2022), while Draft-Sketch-Prove shows LLMs prompted with informal sketches outperform formal-only baselines (Jiang et al., 2023). LeanStar (Lin et al., 2024) interleaves informal reasoning with tactic prediction, typically via supervised fine-tuning on synthetic data. A central open question remains: How do informal reasoning patterns inform formal proving within a model’s internal representations? Our approach explores this via steering vectors in activation space, injecting implicit natural-language “thoughts” at inference to guide proof steps without heavy fine-tuning, probing the latent link between informal intuition and formal reasoning.
Language Model Representation of Concepts Our method is motivated by prior work demonstrating that task or feature concepts can be represented as linear directions in the activation space of LLMs when given appropriate contextual examples. Hendel et al. (2023) and Todd et al. (2024) show that a context (e.g. a prompt) can induce a task embedding in some activation subspace. A broader literature examines linear decompositions of features such as truthfulness (Azaria & Mitchell, 2023; Li et al., 2023; Marks & Tegmark, 2024), sentiment (Tigges et al., 2024), harmlessness (Zou et al., 2025; Zheng et al., 2024), sycophancy or alignment steering (Perez et al., 2023; Panickssery et al., 2024a; Sharma et al., 2024), factual knowledge (Gurnee & Tegmark, 2024), and refusal behavior (Arditi et al., 2024). Relatedly, unsupervised methods like sparse autoencoders have been used to extract concept directions in hidden space (Bricken et al., 2023; Huben et al., 2024; Templeton et al., 2024). These works generally share the hypothesis that LLMs encode high-level features or concepts as (approximately) linear directions in activation space (Elhage et al., 2021; Mikolov et al., 2013; Nanda et al., 2023; Park et al., 2024). While recent work cautions that not all features may admit clean linear representations (Engels et al., 2025), the linearity assumption has proven effective in practice for concept erasure, model steering, and interpretability (Beaglehole et al., 2025; Zhao et al., 2025a; Shah et al., 2025). Within this framework, we explore whether generating a natural-language informal explanation can itself be represented as a linear direction, and how injecting this direction at inference improves formal theorem proving performance.
Mechanistic Interpretability and Activation Patching A rich body of work in mechanistic interpretability seeks to locate, analyze, and manipulate internal representations in transformer-based models. Early studies localized factual knowledge or associative memory to particular neurons or circuits (Meng et al., 2022), and probed hidden layers for high-level features (Li et al., 2024; Dong et al., 2023). The idea of implicit evaluation i.e. measuring latent capability of a model rather than just output behavior, has been developed to complement benchmark-based evaluation (Dong et al., 2023). One particularly relevant tool is activation patching (or residual stream intervention) (Vig et al., 2020; Variengien & Winsor, 2023), wherein one alters specific activations in a layer (often in the residual stream) at inference time to influence model outputs. Recent works have used this to edit factual associations or steer behavior (Zhao et al., 2025b). The residual stream is a high-dimensional accumulator of intermediate features (propagated via skip connections) that each transformer layer refines or adds to; intervening there offers a principled way to steer model behavior (Elhage et al., 2021). In the domain of theorem proving, activation interventions provide a promising lens into how a model processes and integrates informal guidance with formal reasoning. In our approach, we generate steering vectors (patches) that, when applied to the residual stream at certain layers, help the model emit an internal “natural-language thought” prior to or alongside tactic prediction. We find that effective steering tends to realign activations such that the output of the model conforms to a structured reasoning format, thereby improving downstream proof search and reducing failure modes.
Building on these insights, we adapt activation interventions to theorem proving by treating informal reasoning as a direction in activation space. We then show how extracting and injecting this direction can guide proof generation. The next section details the architecture, design choices, and evaluation of this steering approach.
3 Steering for improving theorem proving
We aim to steer theorem-proving models toward using informal reasoning via activation interventions. This section describes (i) how we compute steering vectors, (ii) how we choose layers and apply steering, and (iii) efficiency and practical considerations.
3.1 Intuition and Overview
Modern transformer models often encode high-level semantic behavior (e.g. reasoning, style) as approximately linear directions in activation space (residual streams) (Zou et al., 2025; Elhage et al., 2021). We exploit this property: given pairs of prompts that differ only by the presence of natural-language reasoning (but otherwise describe the same proof step), we can estimate a vector in activation space that points in the “informal reasoning” direction. At inference time, adding this vector (appropriately scaled) nudges the model toward producing reasoning-augmented proofs.
In the rest of this section, we detail how we compute these steering vectors, how we pick which layers to intervene on, and how we incorporate them efficiently at inference.
3.2 Constructing Steering Vectors
We adopt a difference-of-means (contrastive) method (Belrose et al., 2023), which has been effective at extracting feature directions across multiple domains (e.g. refusal, truthfulness) (Arditi et al., 2024; Panickssery et al., 2024b; Marks & Tegmark, 2024). Let $\mathcal{D}=\{(p_{i},p_{i}^{+})\}_{i=1}^{N}$ be a dataset of paired prompts, where $p_{i}$ is a standard formal proof prompt and $p_{i}^{+}$ is its version augmented with explicit natural-language reasoning. We feed both prompts into our model $\mathcal{M}$ and extract the residual stream activations at a chosen layer $\ell$. Denote the activation at the final token position by $\mathbf{v}_{\ell}^{p}=\text{resid}(\mathcal{M}(p),\ell)$.
Then the steering vector $\mathbf{u}^{(\ell)}$ is computed as:
$\mathbf{u}^{(\ell)}=\frac{1}{|\mathcal{D}^{+}|}\sum_{p^{+}\in\mathcal{D}^{+}}\mathbf{v}_{\ell}^{p^{+}}-\frac{1}{|\mathcal{D}^{-}|}\sum_{p^{-}\in\mathcal{D}^{-}}\mathbf{v}_{\ell}^{p^{-}}$ (1)
Here $\mathcal{D}^{+}$ and $\mathcal{D}^{-}$ are the two halves of the prompt-pair dataset (augmented and unaugmented), so the subtraction isolates the direction most correlated with natural-language reasoning while largely cancelling out shared biases or irrelevant activation patterns.
Figure 1: Steering Vectors are computed as difference of means for $p^{+}$ and $p$
We emphasize that this method requires only forward passes; no gradient-based finetuning or parameter updates; making it computationally lightweight.
3.3 Layer Selection and Intervention Strategy
Not all layers are equally responsive to steering, so we first perform an activation analysis to find suitable intervention points. For each layer $\ell$, we measure:
$\text{sim}(\ell)=\frac{1}{|\mathcal{D}|}\sum_{(p,p^{+})\in\mathcal{D}}\cos(\mathbf{v}_{\ell}^{p},\mathbf{v}_{\ell}^{p^{+}})$ (2)
Empirically, we observe that in early layers, $\text{sim}(\ell)$ is close to 1 (i.e. little divergence), but in deeper layers it drops and exhibits local minima (“valleys”) as seen in Figure 2.
Figure 2: Cosine similarity across model’s residual stream activations for each layer.
We find that in early layers, activations remain highly similar between $p$ and $p^{+}$, but as we go deeper, the cosine similarity drops and exhibits local minima. Intuitively, these are the layers where the representations diverge most when informal reasoning is introduced (see Figure 2). Thus, these “valley” layers are promising candidates for steering. These valleys likely correspond to points where the internal representation of the model is most sensitive to reasoning-specific perturbations.
We select intervention layers where (a) cosine similarity dips significantly, and (b) representation is still semantically stable. At inference, we inject the steering vector into those layers:
$\mathbf{v}_{\ell}^{\prime}=\mathbf{v}_{\ell}+\alpha\cdot\mathbf{u}^{(\ell)}$ (3)
We treat $\alpha$ as a hyperparameter and specify that it will be tuned on held-out validation data. The method instantiates a sweep to find the optimal balance between influencing reasoning and maintaining validity. We place constraints (e.g. an upper bound on $\alpha$) to avoid destabilizing proofs. We present the complete pseudocode of our approach in Algorithm 1.
Algorithm 1 Activation Steering for Theorem Proving
0: Contrastive prompt pairs $D=\{(p_{i},p_{i}^{+})\}_{i=1}^{N}$ where $p_{i}$ are standard prompts and $p_{i}^{+}$ include natural language reasoning
0: Language model $M$ with $L$ layers
0: Test theorem $T$ to prove
0: Enhanced proof generation
1:Offline: Create Steering Vectors
2:for each layer $\ell\in\{1,\ldots,L\}$ do
3:  $\bar{v}^{+}\leftarrow\frac{1}{N}\sum_{i=1}^{N}\text{resid}(M(p_{i}^{+}),\ell)$ {Mean activation for augmented prompts}
4:  $\bar{v}^{-}\leftarrow\frac{1}{N}\sum_{i=1}^{N}\text{resid}(M(p_{i}),\ell)$ {Mean activation for standard prompts}
5:  $u^{(\ell)}\leftarrow\bar{v}^{+}-\bar{v}^{-}$ {Steering vector at layer $\ell$}
6:end for
[…]
8:Online: Apply Steering During Inference
9: Initialize proof context with theorem $T$
10:while proof incomplete and within computational budget do
11:  Perform forward pass through model
12:  for each intervention layer $\ell\in\mathcal{L}$ do
13:   $v^{\prime}_{\ell}\leftarrow v_{\ell}+\alpha\cdot u^{(\ell)}$ {Inject steering vector}
14:  end for
15:  Generate next proof step using modified activations
16:  Validate and apply proof step if correct
17:end while
18:return Generated proof or failure
4 Experimental Setup
4.1 Models
We conduct experiments on three open-source 7B-parameter language models optimized for mathematical reasoning:
  • Llemma-7B (Azerbayev et al., 2024): Built on Code Llama architecture, pre-trained on mathematical texts and formal mathematics corpora.
  • InternLM2-7B (Ying et al., 2024): LLaMA-based architecture with extended training on mathematical problem-solving.
  • InternLM2.5-StepProver (Wu et al., 2024): Enhanced variant with expert iteration on large-scale Lean problems.
All models share transformer architectures with consistent residual stream dimensionality, facilitating uniform steering vector extraction.
4.2 Data and Vector Construction
We construct steering vectors from the Lean-STaR dataset(Lin et al., 2024), randomly sampling 10,000 theorem-proof pairs. Each datapoint $(p,p^{+})$ consists of a formal proof step $p$ and its augmented version $p^{+}$ containing explicit natural language reasoning.
For robustness, we generate model responses $r_{p}$ and $r_{p^{+}}$ for each prompt pair and retain only instances where both responses are valid proof steps with $r_{p}\neq r_{p^{+}}$. This filtering yields approximately 7,400 high-quality contrastive pairs.
Vector construction requires only forward passes through the model, consuming $\sim 1$ GPU-hour on NVIDIA A100, a 60× reduction compared to LoRA fine-tuning while achieving superior performance gains.
4.3 Chosen Layers and $\alpha$
Based on activation similarity analysis (Section 3.3), we apply steering vectors to Layers 22, 25 in InternLM2-7B and Layers 24, 25 in Llemma-7B. These correspond to layers where the model’s representation diverges most from baseline while maintaining semantic coherence. We empirically determine $\alpha=0.8$ through validation experiments.
4.4 Evaluation Benchmarks
miniF2F Our primary evaluation utilizes miniF2F (Zheng et al., 2022), a standardized benchmark comprising 244 theorems drawn from mathematical competitions (AMC, AIME, IMO). These problems span algebra, number theory, geometry, and combinatorics with varying proof complexities. We employ two proof search strategies:
  • Best-first search: Expansion budget $N\in\{50,600\}$, tactic sampling width $S=32$, selecting states by cumulative log-probability
  • Parallel sampling: $K$ independent proof attempts to accommodate probability shifts from natural language injection
PutnamBench We additionally evaluate on PutnamBench (Tsoukalas et al., 2024), containing problems from the William Lowell Putnam Mathematical Competition. Following standard evaluation protocol, we test on both the Lean subset (657 problems) and Rocq subset (412 problems) using the benchmark’s default search parameters: $N=600$ expansion budget with $S=32$ tactic width for best-first search, and parallel sampling with $K=2$ attempts. This benchmark features more challenging problems with longer average proof lengths, providing a stringent test of steering effectiveness on complex mathematical reasoning.
Evaluation Metrics.  Following established practice in neural theorem proving, we report pass rates (percentage of theorems successfully proved within computational budget) as our primary metric. For detailed analysis, we additionally examine proof characteristics including average length, tactic distribution, and the frequency of intermediate lemma usage (via the have tactic) to understand how steering affects proof structure and strategy.
5 Results and Analysis
Our experiments were designed to address the following research questions:
@@ 5 Results and Analysis
    Does activation steering improve theorem-proving performance of existing models?
@@
    Which layers contribute most effectively when steered?